OpenStudy (anonymous):

Please help! Show steps please, thanks!! ;) If Zachary earns $90 simple Interest on $750 at 6% for 2 years, what is his new saving account balance? @Mathmatics

OpenStudy (anonymous):

OH spelled Mathematics wrong

OpenStudy (anonymous):

you know some formulas to solve this problem dont you?

OpenStudy (anonymous):

yeah but I forgot.... :\

OpenStudy (anonymous):

Consider this one:A=P+Prt, where P is the principal amount and r is the rate and t is the time

OpenStudy (anonymous):

A=750(1+(.06)(2)) =750(1+.12) =750(1.12) =840

OpenStudy (anonymous):

or just by common sense, if he had 750 in his account and he earned 90 dollars, then he had 750+90=840 dollars in his account

OpenStudy (anonymous):

OH ook, Thankss!!!! ;]